# Holographic Screen Sequestration

### 11 Dec 2023 v1

@@ III.1.3 Inner wedge $I_{W}[X]$
In the inner wedge, screens with early timelike flows must asymptote to $N_{+\ell}[X]$ by sequestration. These screens can again turn spacelike outwards, but like in the future wedge, will inevitably run into $X$. Again, this case can be ruled out by the presence of forbidden loops as shown in Fig. 4, so screens with early timelike flows can only have late timelike flows that asymptote to $N_{-k}[X]$ by sequestration. Screens with early spacelike flows can have spacelike segments either to the future or past of the Cauchy surface $\Sigma$ like screens in the outer wedge, but must turn towards the past in a timelike fashion by virtue of the area law and again asymptote to $N_{-k}[X]$ for late flows.
@@ III.1.4 Causal past $I^{-}[X]$
Similar to screens with early spacelike flows in the outer wedge, screens with early timelike flows in the causal past of $X$ can only start at the HRT surface $X$ due to sequestration. From here, the screen can have late timelike flows toward past infinity, or turn spacelike outwards and asymptote to $N_{-\ell}[X]$ for late spacelike flows. Screens with early spacelike flows in the past of $X$ must asymptote, backwards along the flow, to $N_{-k}[X]$ by sequestration, then continue to have a late timelike flow to past infinity following the area law, or asymptote to $N_{-\ell}[X]$ with a late spacelike flow by sequestration.
@@ III.2 Nonsymmetric spacetimes
As before, let $\sigma(\tau)$ denote leaves foliating a holographic screen $H$ with $\tau$ parameterizing the flow along the screen so that $A[\sigma(\tau_{1})]<A[\sigma(\tau_{2})]$ for $\tau_{1}<\tau_{2}$. We will now consider the case in which the null congruence $N_{\pm u}[X]$ intersects $H$ in an oblique manner such that the intersection $H\cap N_{\pm u}[X]=Y$ is some compact spacelike surface that is not a leaf, i.e., not a marginal surface. More generally, since $k,\ell$ are defined everywhere by the outgoing null congruences shot out from the leaves $\sigma(\tau)$ of $H$, and parallel transporting along themselves, the null congruences fired from $X$ in nonsymmetric spacetimes will no longer be generated by $k,\ell$ as they are not necessarily orthogonal to $X$ [^4] See the discussion in Step 2, Sec. 5.2 of Ref. Engelhardt:2018kcs , for example. With this caveat in mind, we will call the $X$-orthogonal null congruences $N_{\pm u}[X]$ for $u=\tilde{k},\tilde{\ell}$, which are in general nonparallel to the $\sigma(\tau)$-orthogonal $k,\ell$ unless the spacetime exhibits sufficient (e.g., spherical) symmetry. The geometry of these surfaces is illustrated in Fig. 6 for the case of $Y=H\cap N_{\tilde{k}}[X]$ without loss of generality (a similar cartoon can be drawn for $Y=H\cap N_{\tilde{\ell}}[X]$).
@@ III.2.2 Local area law for infinitesimal elements of proper area
for some scalar fields $\alpha$ and $\beta$. By Theorem IV.2 of Ref. Bousso:2015qqa , $\alpha<0$ everywhere on a future holographic screen $H$. We are still free to normalize the null generators $k^{\mu}$ and $\ell^{\mu}$ such that $k\cdot\ell=-1$ and $\ell^{\mu}$ is an affine generator. To preserve the cross-normalization however, rescaling $\ell^{\mu}$ also means rescaling $k^{\mu}$. Under this rescaling, $k^{\mu}$ will generically not be affinely parameterized, but will remain proportional to an affine generator, i.e.,
@@
where $c_{k}$ is a scalar. Since $\ell^{\mu}$ affinely parameterizes null congruences $N_{\ell}[\sigma]$ fired from leaves of the screen, the fractional rate of change of cross-sectional area of $N_{\ell}[\sigma]$ is given precisely by the null expansion
@@
Eq. (13) then follows from the fact that $\theta^{\sigma(\tau)}_{k}=0$ everywhere on $\sigma(\tau)$. The monotonicity of this infinitesimal area law follows from the property that $\alpha(p)<0$ and $\theta^{\sigma(\tau)}_{\ell}<0$ at every point $p\in\sigma(\tau)$. ∎
@@
Eq. (13) also holds for past holographic screens.
@@ III.3.3 $N_{\pm\tilde{k}}[X]$ cannot contain a complete slice of $H$
On the other hand, for a screen sequestered to the exterior of $X$ to simultaneously intersect $N_{+\tilde{k}}[X]$ at $Y$ and respect monotonicity of area along the screen, the screen must have a semi-infinite null segment with area flow from future null infinity to $Y$ along $N_{+\tilde{k}}[X]$, then escape from $Y$ to the outer wedge of $X$. In the same vein, a screen sequestered to the past of $X$ will necessarily have a null segment with flow from past null infinity to $Y$ along $N_{-\tilde{k}}[X]$, and leave $Y$ to the past of $X$.
@@
No forbidden loops exist for such outgoing screens from $N_{\pm\tilde{k}}[X]$, so these constructions (shown in Fig. 10) are in principle allowed. While these are still holographic screens in a general sense, the presence of a semi-infinite null segment means that they run afoul of a genericity condition in Definition II.8 of Ref. Bousso:2015qqa that defines regular holographic screens. As reviewed in Sec. II, this condition requires a regular holographic screen to have no set of leaves of nonzero measure on which the tangent to the screen’s flow is null and outward-directed. If we restrict ourselves to only regular holographic screens, then we are forced to forgo the remaining two types of irregular screens containing semi-infinite null segments along $N_{\pm\tilde{k}}[X]$.
@@ V.2 Other extremal surfaces
In proving results about sequestration, we worked exclusively with HRT surfaces, which in particular are extremal surfaces that are maximin. There are of course other types of extremal surfaces: minimax extremal surfaces, which have maximal area on a Cauchy surface and have the least area among the maximal area slices of a complete family of Cauchy surfaces, or even “maximax” surfaces, which are largest area surfaces among such a family of maximal area surfaces. Examples of minimax and maximax extremal surfaces are the equatorial sphere of the throat of a closed bouncing FLRW cosmology and the equatorial sphere of the maximal volume slice of a Big Bang-Big Crunch FLRW cosmology, respectively. In such cases, however, the extremal surface has little sequestering power. In a diagram that is analogous to Fig. 3, the extremal surface $X$ would look like an attractor rather than a saddle. The absence of any lines that represent surfaces of increasing area away from $X$ prevent us from forming the forbidden loops that were necessary to argue for sequestration, save for the loop shown in the bottom right of Fig. 3. We therefore conclude that future (past) holographic screens are only forbidden from crossing from $O_{W}[X]$ into $I^{-}[X]$ ($I^{+}[X]$) in the nonmaximin case.
@@ V.3 Semiclassical generalizations
Acknowledgments: We thank Bartek Czech, Edgar Shaghoulian, and Mark Van Raamsdonk for comments and discussions during the preparation of this manuscript. A.C.-D. was supported for a portion of this work by the Natural Sciences and Engineering Research Council of Canada (NSERC), [funding reference number PDF-545750-2020]. P.L. is supported by the Four Year Doctoral Fellowship at the University of British Columbia. G.N.R. is supported by the James Arthur Postdoctoral Fellowship at New York University. This work has benefited from discussions at the QIMG 2023 long term workshop (YITP-T-23-01) held at the Yukawa Institute for Theoretical Physics, Kyoto University.

### 07 Feb 2024 v2

@@ III.1.3 Inner wedge $I_{W}[X]$
In the inner wedge, screens with early timelike flows must asymptote to $N_{+\ell}[X]$ by sequestration. These screens can again turn spacelike outward, but like in the future wedge, will inevitably run into $X$. Again, this case can be ruled out by the presence of forbidden loops as shown in Fig. 5, so screens with early timelike flows can only have late timelike flows that asymptote to $N_{-k}[X]$ by sequestration. Screens with early spacelike flows can have spacelike segments either to the future or past of the Cauchy surface $\Sigma$ like screens in the outer wedge, but must turn toward the past in a timelike fashion by virtue of the area law and again asymptote to $N_{-k}[X]$ for late flows.
@@ III.1.4 Causal past $I^{-}[X]$
Similar to screens with early spacelike flows in the outer wedge, screens with early timelike flows in the causal past of $X$ can only start at the HRT surface $X$ due to sequestration. From here, the screen can have late timelike flows toward past infinity, or turn spacelike outward and asymptote to $N_{-\ell}[X]$ for late spacelike flows. Screens with early spacelike flows in the past of $X$ must asymptote, backward along the flow, to $N_{-k}[X]$ by sequestration, then continue to have a late timelike flow to past infinity following the area law, or asymptote to $N_{-\ell}[X]$ with a late spacelike flow by sequestration.
@@ III.2 Nonsymmetric spacetimes
As before, let $\sigma(\tau)$ denote leaves foliating a holographic screen $H$ with $\tau$ parametrizing the flow along the screen so that $A[\sigma(\tau_{1})]<A[\sigma(\tau_{2})]$ for $\tau_{1}<\tau_{2}$. We will now consider the case in which the null congruence $N_{\pm u}[X]$ intersects $H$ in an oblique manner such that the intersection $H\cap N_{\pm u}[X]=Y$ is some compact spacelike surface that is not a leaf, i.e., not a marginal surface. More generally, since $k,\ell$ are defined everywhere by the outgoing null congruences shot out from the leaves $\sigma(\tau)$ of $H$, and parallel transporting along themselves, the null congruences fired from $X$ in nonsymmetric spacetimes will no longer be generated by $k,\ell$ as they are not necessarily orthogonal to $X$ [^4] See the discussion in Step 2, Sec. 5.2 of Ref. Engelhardt:2018kcs , for example. With this caveat in mind, we will call the $X$-orthogonal null congruences $N_{\pm u}[X]$ for $u=\tilde{k},\tilde{\ell}$, which are in general nonparallel to the $\sigma(\tau)$-orthogonal $k,\ell$ unless the spacetime exhibits sufficient (e.g., spherical) symmetry. The geometry of these surfaces is illustrated in Fig. 6 for the case of $Y=H\cap N_{\tilde{k}}[X]$ without loss of generality (a similar cartoon can be drawn for $Y=H\cap N_{\tilde{\ell}}[X]$).
@@ III.2.2 Local area law for infinitesimal elements of proper area
for some scalar fields $\alpha$ and $\beta$. By Theorem IV.2 of Ref. Bousso:2015qqa , $\alpha<0$ everywhere on a future holographic screen $H$. We are still free to normalize the null generators $k^{\mu}$ and $\ell^{\mu}$ such that $k\cdot\ell=-1$ and $\ell^{\mu}$ is an affine generator. To preserve the cross-normalization however, rescaling $\ell^{\mu}$ also means rescaling $k^{\mu}$. Under this rescaling, $k^{\mu}$ will generically not be affinely parametrized, but will remain proportional to an affine generator, i.e.,
@@
where $c_{k}$ is a scalar. Since $\ell^{\mu}$ affinely parametrizes null congruences $N_{\ell}[\sigma]$ fired from leaves of the screen, the fractional rate of change of cross-sectional area of $N_{\ell}[\sigma]$ is given precisely by the null expansion
@@
Equation (13) then follows from the fact that $\theta^{\sigma(\tau)}_{k}=0$ everywhere on $\sigma(\tau)$. The monotonicity of this infinitesimal area law follows from the property that $\alpha(p)<0$ and $\theta^{\sigma(\tau)}_{\ell}<0$ at every point $p\in\sigma(\tau)$. ∎
@@
Equation (13) also holds for past holographic screens.
@@ III.3.3 $N_{\pm\tilde{k}}[X]$ cannot contain a complete slice of $H$
On the other hand, for a screen sequestered to the exterior of $X$ to simultaneously intersect $N_{+\tilde{k}}[X]$ at $Y$ and respect monotonicity of area along the screen, the screen must have a semi-infinite null segment with area flow from future null infinity to $Y$ along $N_{+\tilde{k}}[X]$, then escape from $Y$ to the outer wedge of $X$. In the same vein, a screen sequestered to the past of $X$ will necessarily have a null segment with flow from past null infinity to $Y$ along $N_{-\tilde{k}}[X]$, and leave $Y$ to the past of $X$. No forbidden loops exist for such outgoing screens from $N_{\pm\tilde{k}}[X]$, so these constructions (shown in Fig. 10) are in principle allowed. While these are still holographic screens in a general sense, the presence of a semi-infinite null segment means that they run afoul of a genericity condition in Definition II.8 of Ref. Bousso:2015qqa that defines regular holographic screens. As reviewed in Sec. II, this condition requires a regular holographic screen to have no set of leaves of nonzero measure on which the tangent to the screen’s flow is null and outward-directed. If we restrict ourselves to only regular holographic screens, then we are forced to forgo the remaining two types of irregular screens containing semi-infinite null segments along $N_{\pm\tilde{k}}[X]$.
@@ V.2 Other extremal surfaces
In proving results about sequestration, we worked exclusively with HRT surfaces, which in particular are extremal surfaces that are maximin. There are of course other types of extremal surfaces: minimax extremal surfaces, which have maximal area on a Cauchy surface and have the least area among the maximal area slices of a complete family of Cauchy surfaces, or even “maximax” surfaces, which are largest area surfaces among such a family of maximal area surfaces. Examples of minimax and maximax extremal surfaces are the equatorial sphere of the throat of a closed bouncing Friedmann–Lemaître–Robertson–Walker (FLRW) cosmology and the equatorial sphere of the maximal volume slice of a big bang–big crunch FLRW cosmology, respectively. In such cases, however, the extremal surface has little sequestering power. In a diagram that is analogous to Fig. 3, the extremal surface $X$ would look like an attractor rather than a saddle. The absence of any lines that represent surfaces of increasing area away from $X$ prevent us from forming the forbidden loops that were necessary to argue for sequestration, save for the loop shown in the bottom right of Fig. 3. We therefore conclude that future (past) holographic screens are only forbidden from crossing from $O_{W}[X]$ into $I^{-}[X]$ ($I^{+}[X]$) in the nonmaximin case.
@@ V.3 Semiclassical generalizations
Acknowledgments: We thank Bartek Czech, Edgar Shaghoulian, and Mark Van Raamsdonk for comments and discussions during the preparation of this manuscript. A. C.-D. was supported for a portion of this work by the Natural Sciences and Engineering Research Council of Canada (NSERC), [funding reference number PDF-545750-2020], and by Grant #62312 from the John Templeton Foundation for “The Quantum Information Structure of Spacetime” (QISS) project. P.L. is supported by the Four Year Doctoral Fellowship at the University of British Columbia. G.N.R. is supported by the James Arthur Postdoctoral Fellowship at New York University. This work has benefited from discussions at the QIMG 2023 long term workshop (YITP-T-23-01) held at the Yukawa Institute for Theoretical Physics, Kyoto University.